OpenStudy (anonymous):

What are the last 2 digits of the number 43^34?

OpenStudy (unklerhaukus):

what are the first few powers of 43,

Parth (parthkohli):

Mathematica: Table[43^n, {n, 1, 5, 1}] {43, 1849, 79507, 3418801, 147008443}

Parth (parthkohli):

Looks like it returns to 43 every 5 powers.

OpenStudy (anonymous):

\[43^{34}= 43*43*43^{32}=1849*43^{32}=1849*43*43^{31}=147008443*43^{31}\]

OpenStudy (anonymous):

for every power of 5 we have a 43 and since its a 3 at one's place we can have 9 at the one place of the 34th power

OpenStudy (phi):

use modulo 100 on the powers of 43 43^0 01 43^1 43 43^2 49 mod 100 43^3 07 43^4 01 it starts repeating 43^32 = 43^0 mod 100 43^34= 43^2 =49 as the last 2 digits